Consider the reaction between NO and Cl2 to form NOCl: 2NO(g)+Cl2(g)⇌2NOCl(g) A reaction mixture at a certain temperature initially contains only [NO]= 0.67 M and [Cl2]= 0.65. After the reaction comes to equilibrium, the concentration of NOCl is 0.15 M. Find the value of the equilibrium constant (Kc) at this temperature. Express your answer using two significant figures.
Answer #1

2 NO(g) +   Cl2(g)   ⇌   2 NOCl(g)

0.67           0.65                 0

0.67-2x    0.65-x                2x

At equilibrium :

concentration of NOCl = 2x = 0.15 M

x = 0.075 M

[NO] = 0.67 - 2x 0.075 = 0.52 M

[Cl2] = 0.65 - 0.075 = 0.575 M

Kc = [NOCl]^2 / [NO]^2[Cl2]

    = (0.15)^2 / (0.52)^2(0.575)

Kc = 0.14

equilibrium constant = 0.14